# Mixmaster chaos in an AdS black hole interior

### v1

@@ 8 Discussion
In order to obtain an explicit realisation of mixmaster chaos in an asymptotically AdS setting we have considered a cohomogeneity-one Ansatz (i.e. the fields only depend on one coordinate) with a diagonal metric. This Ansatz consistently leads to solutions to the full equations of motion. However, clearly, it does not give a fully generic approach to the interior singularity. A generic solution will be inhomogeneous and will develop off-diagonal components in the metric.
@@
[…]
@@ Appendix A Equations of motion
Figure 17: Irrelevance of the mass and cosmological constant terms towards the interior. The first three curves show that fluxes in the vector field equations (152), (153) and (154) become conserved towards the interior. The final curve shows that the first term in (149) equals the final term at late interior times. The collapse of the Einstein-Rosen bridge is shown as a vertical dashed line for reference. Numerical parameters are as in Fig. 2.
@@ Appendix D Hecke operators and automorphic waveforms
[…]
@@
[…]

### v4

@@ 8 Discussion
In order to obtain an explicit realisation of mixmaster chaos in an asymptotically AdS setting we have considered a cohomogeneity-one Ansatz (i.e. the fields only depend on one coordinate) with a diagonal metric and no longitudinal component to the vector fields. This Ansatz consistently leads to solutions to the full equations of motion. However, clearly, it does not give a fully generic approach to the interior singularity. A generic solution will be inhomogeneous and will develop off-diagonal components in the metric.
@@
The dynamics of the large and small triangles have been further studied in [87, 88, 89].
@@
$S({\varepsilon})=\frac{V}{(\ell_{\text{Pl}})^{3}}\log\frac{{\varepsilon}}{12}\,.$ (148)
Here $V$ is spatial volume of the interior slice, in units of a ‘Planck scale’ cutoff $\ell_{\text{pl}}$. Even with the entropy (148) regulated at short distances, it still suffers from a long distance divergence because the volume is infinite. This large number of states is in tension with the finite black hole entropy, and is a version of the ‘bags of gold’ paradox for eternal AdS black holes [90].
@@ Appendix A Equations of motion
Figure 17: Irrelevance of the mass and cosmological constant terms towards the interior. The first three curves show that fluxes in the vector field equations (152), (153) and (154) become conserved towards the interior. The final curve shows that the first term in (149) equals the final term at late interior times (this curve was made by using (149) to obtain an expression for $e^{H}(e^{-H}F/z^{3})^{\prime}$ and plotting that). The collapse of the Einstein-Rosen bridge is shown as a vertical dashed line for reference. Numerical parameters are as in Fig. 2.
@@ Appendix D Hecke operators and automorphic waveforms
$\gamma=\begin{pmatrix}a_{1}&b_{1}\\ c_{1}&d_{1}\end{pmatrix}\qquad\text{and}\qquad m_{p}=\begin{pmatrix}a&b\\ c&d\end{pmatrix},$ (169)
@@
$\alpha_{p}=\begin{pmatrix}a&2b\\ 0&d\end{pmatrix},\qquad\alpha_{p}^{\prime}=\begin{pmatrix}a^{\prime}&2b^{\prime}\\ 0&d^{\prime}\end{pmatrix},\qquad\text{and}\qquad\gamma=\begin{pmatrix}a_{1}&2b_{1}\\ 2c_{1}&d_{1}\end{pmatrix}\,.$ (173)